OpenStudy (winsteria):

Which answer describes the solution to each equation? x-6/2 = x + 13 A. one solution B. no solution C. an infinite number of solutions

OpenStudy (winsteria):

@KamiBug @acxbox22

OpenStudy (alexandervonhumboldt2):

x-x=13+6/2 0=16 false sono solution

OpenStudy (winsteria):

what @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

??????????

OpenStudy (alexandervonhumboldt2):

B

OpenStudy (winsteria):

oh okay thank you sorry it was glitching and it showed up as question marks

OpenStudy (alexandervonhumboldt2):

ok